# Let it Snow! Animating 3D Gaussian Scenes with Dynamic Weather Effects via Physics-Guided Score Distillation

### v2

@@ 3 Method
In this section we introduce our Physics-Guided Score Distillation framework for editing 3D Gaussian Splatting scenes scene-wide, multi-particle edits. As illustrated in Fig. 2, our approach consists of two components: First, we use Material Point Method (MPM) simulation to generate physically plausible reference trajectories for dynamic particles (Sec. 3.2). These trajectories serve as a guidance prior rather than fixed motion. Second, we learn a recurrent neural dynamics model that predicts dynamic Gaussian evolution over time, optimized through Physics-Guided Score Distillation (Sec. 3.3), a joint loss combining Video Score Distillation Sampling (Video-SDS) for photorealistic refinement of both motion and appearance with physics regularization for motion plausibility. We begin by providing a brief background about the foundational techniques (Sec. 3.1) that form the basis of our approach. See Appendix B.1 for implementation details.
@@ 3.2 Physics-Based Motion Prior
Due to the coarse grid resolution of MPM, we refine particle-scene interactions through mesh-based collision handling. For snow, we project particles onto mesh surfaces and interpolate with nearby surface Gaussians to create natural accumulation. For rain, we implement a 3D wetness grid that tracks moisture with Gaussian smoothing and temporal decay. For sandstorms, we displace particles along surface normals with anisotropic scaling to create thin accumulations. When particles become inactive ($a_{i}^{t}=1$ to $a_{i}^{t+1}=0$), their final positions and rotations are recorded as collision states. These simulated trajectories provide physically plausible reference motion that guides our neural dynamics model. See Appendix B.2 and B.3 for additional collision handling and simulation details.
@@ 3.3 Physics-Guided Score Distillation
Rather than treating physics simulation as fixed motion, we use it as a guidance prior to optimize a neural dynamics model through Video-SDS. For each dynamic particle from the MPM simulation, we create a corresponding dynamic Gaussian kernel $\mathcal{G}_{g}(t)=\{\mathbf{x}_{g}^{\text{init}}(t),\mathbf{R}_{g}^{\text{init}},\mathbf{A}_{g}^{\text{active}},\mathbf{A}_{g}^{\text{collided}},\mathbf{v}_{g}^{\text{init}}(t)\}$, where $\mathbf{v}_{g}^{\text{init}}(t)$ is the velocity of each gaussian that is provided along the position $\mathbf{x}_{g}^{\text{init}}(t)$ directly from the motion simulation. Each Gaussian also maintains state-dependent appearance parameters based on its physical state determined by the physics simulation: $\mathcal{A}_{g}^{\text{active}}=\{sigma_{g}^{\text{active}},\mathbf{S}_{g}^{\text{active}},\mathbf{C}_{g}^{\text{active}}\}$ for actively moving Gaussians and $\mathcal{A}_{g}^{\text{collided}}$ for Gaussians that have accumulated on static geometry. For notational clarity, we omit the explicit state subscripts and use $\mathcal{A}_{g}=\{\sigma_{g}^{\text{collided}},\mathbf{S}_{g}^{\text{collided}},\mathbf{C}_{g}^{\text{collided}}\}$ to refer to the appropriate appearance parameters based on the Gaussian’s current physical state. Appearance parameters are initialized using a large language model (LLM) that generates appropriate baseline values $\mathcal{A}_{g}^{\text{init}}$ from text prompts describing the desired weather effect. See Appendix B.4 for details on LLM initialization.
Recurrent Neural Dynamics Model. To model the continuous evolution of the dynamic Gaussians, we learn a recurrent neural dynamics model that predicts dynamic Gaussian evolution over time. At each timestep $t$, the model takes as input: (1) the previous rendered state $\{\mathbf{x}_{g}(t-1),\mathbf{R}_{g}(t-1),\mathcal{A}_{g}(t-1)\}$ from the previous timestep, where $\mathbf{x}_{g}(t-1)$ is the position of Gaussian $g$ at time t1t-1, $\mathbf{R}_{g}(t-1)$ is the rotation of Gaussian $g$ and $\mathcal{A}_{g}$ represents the appearance parameters of Gaussian $g$ (opacity, scale, and color), (2) the physics velocity $\mathbf{v}_{g}^{\text{init}}(t)$ of Gaussian $g$ at time $t$ from simulation, and (3) the current timestep $t$. We implement separate MLPs for each physical state, with each MLP processing Gaussians independently. The positions, velocities, and timestep are processed through learned embedders (Fourier feature encoding for spatial inputs, sinusoidal encoding for time) before being concatenated with the current appearance parameters $\mathcal{A}_{g}(t-1)$ and fed to the appropriate state-specific MLP. The recurrence is initialized at $t=0$ using physics simulation positions $\mathbf{x}_{g}^{\text{init}}(0)$ and the LLM-initialized appearance parameters $\mathcal{A}_{g}^{\text{init}}$. See Appendix B.1 for additional details on the network.
@@
Each regularization term has independent weight parameters to provide control over physics-appearance trade-off. See Appendix B.1 for additional details.
@@
Figure 3: Dynamic weather effects created with our physics-guided SDS framework across different scenes (input frames on left). Each column represents an increasing timestep, demonstrating temporal evolution with continuous particle emission and realistic scene interactions. Our approach enables scene-wide modifications where emitted particles interact naturally with static geometry and evolve coherently over time. The snowfall effect is presented in Fig. 1. As the rainfall effect is challenging to visualize, we provide zoomed-in regions to better illustrate the effect. Complete temporal visualizations for all effects are available in our project webpage.
@@ 4.1 Qualitative Results
Fig. 1 (snowfall) and Fig. 3 (all other dynamic weather effects) demonstrate our method’s ability to create visually convincing and temporally consistent dynamic scene modifications across different environments. Our physics-guided score distillation framework enables effects that exhibit physically plausible motion while maintaining photorealistic appearance: snowflakes gently accumulate on surfaces through continuous particle emission, raindrops create wetness patterns, fog diffuses through the scene, and sandstorms envelope objects with physically-driven particle motion. Additionally, Fig. 4 showcases creative appearance variations generated by our framework through text guidance (e.g., purple snow) while maintaining physically plausible motion through our physics-guided optimization. Please refer to our project webpage for additional visualizations across different scenes, effects, and viewpoints.
@@ 4.2 Comparison To Baselines
Table 1: Quantitative Evaluation. We compare against the static 3D editing methods ClimateNeRF [16] and GaussCtrl [33]. Results are shown for different edit categories, with F+S representing results for fog and snow edits (the only effects supported by ClimateNeRF) and All representing results across all edit types.
@@
Baselines. Our work addresses the challenge of introducing dynamic weather effects to static scenes, which differs from existing approaches. Recent 4D editing methods [20, 14] leverage score distillation for scene-wide modifications, however these methods operate in a different setting, applying edits to a fixed set of Gaussians. Similarly, 4D generation techniques like DreamGaussian4D [24] and Animate124 [38] primarily target object-centric animation rather than scene-wide effects. Physics-based methods such as PhysGaussian [34] and Feature Splatting [22] focus on manipulating existing scene elements through simulation but are not designed to synthesize new dynamic elements with photorealistic appearance. RainyGS [3] addresses rain synthesis through physics-based particle systems with rain-specific simulation and rasterization techniques, However it does not generalize to other weather effects. Additionally, we cannot provide direct comparisons as their code is not publicly available. Given these limitations, we compare our approach against static 3D editing methods that represent the most directly comparable baselines. Specifically, we evaluate against ClimateNeRF [16], which adds static weather effects to NeRFs, and GaussCtrl [33], a method for text-guided editing of 3D Gaussian Splatting scenes.
In the appendix, we compare our method against Instruct-4DGS [14], a 4D editing baseline. See Appendix A for the 4D editing baseline comparison details and results.
@@
Comparisons. A qualitative comparison to static 3D editing methods is shown in Fig. 5. We show our results at two different timesteps: one at an early stage of the simulation (Small $t$) and one at the end (Large $t$). In the snowfall example in the first row, the static methods cannot capture the temporal dynamics of continuous particle emission and accumulation, yielding only surface modifications (ClimateNeRF) or appearance changes (GaussCtrl). In the second row, for the fog example, the effect is either unnoticeable (ClimateNeRF) or unrealistic (GaussCtrl). Our physics-guided framework produces effects that exhibit both physically plausible temporal evolution and photorealistic appearance. See Appendix C for additional comparisons across different scenes, effects, and viewpoints.
In Tab. 1 we show a quantitative comparison to these static editing methods. We create moving camera videos from the static method results to enable fair comparison using video-based metrics. ClimateNeRF only performs modifications for the fog and snow effects, thus we show results over these effects (F+S) and also over all weather effects. As illustrated in the table, our method achieves higher scores across both video and image-based metrics, demonstrating that our approach produces high-quality effects even in static frames, while the temporal dimension, enabled by continuous particle emission and physics-guided optimization, further distinguishes our method.
@@ 4.3 Ablations
Fig. 6 and Tab. 2 illustrate that each ablated component leads to a significant degradation in quality. The w/o CH ablation demonstrates that without proper collision handling, dynamic Gaussians fail to interact convincingly with static scene geometry, with snowflakes floating unnaturally above surfaces instead of accumulating realistically. While the w/o App variant achieves physically plausible motion, it produces suboptimal visual appearance due to simplified LLM-initialized parameters. The w/o Motion variant produces incoherent and noisy results, producing significantly lower results than other variants. These results demonstrate that video diffusion models alone cannot reliably generate physically plausible motion for complex multi-particle scenarios and are not designed to introduce new Gaussians. The w/o PG variant demonstrates that fixed motion prevents effective Video-SDS optimization. Without the joint optimization enabled by our physics-guidance prior, the model cannot learn the motion refinements required for the diffusion model to satisfy the text prompt.
In the appendix, we conduct additional experiments, ablating our SDS-adaptive weighting mechanism, our regularization losses and studying the threshold behavior of trajectory drift. See Appendix A for the additional ablations and results.
@@ 5 Conclusion
Appendix
Appendix A Additional Ablations and Comparisons
@@
As shown in Fig. 7 and Tab. 3, Instruct-4DGS produces incoherent results with physically implausible motion patterns. Without a physics-based motion prior to guide the optimization, the method struggles to learn realistic dynamics for complex multi-particle weather phenomena, resulting in unnatural motion and visual artifacts.
@@
Table 3: 4D Editing Baseline Comparison. Quantitative comparison against Instruct-4DGS [14], a recent 4D editing method. Without physics-based motion guidance, Instruct-4DGS struggles to generate coherent multi-particle dynamics, resulting in lower scores across all metrics.
Instruct-4DGS 0.21 0.07 0.57 0.13 0.19
Ours $\mathbf{0.28}$ $\mathbf{0.11}$ $\mathbf{0.89}$ $\mathbf{0.19}$ $\mathbf{0.41}$
@@
Appendix B Additional Details
B.1 Implementation Details

### v4

@@ 3 Method
In this section we introduce our Physics-Guided Score Distillation framework for editing 3D Gaussian Splatting scenes scene-wide, multi-particle edits. As illustrated in Fig. 2, our approach consists of two components: First, we use Material Point Method (MPM) simulation to generate physically plausible reference trajectories for dynamic particles (Sec. 3.2). These trajectories serve as a guidance prior rather than fixed motion. Second, we learn a recurrent neural dynamics model that predicts dynamic Gaussian evolution over time, optimized through Physics-Guided Score Distillation (Sec. 3.3), a joint loss combining Video Score Distillation Sampling (Video-SDS) for photorealistic refinement of both motion and appearance with physics regularization for motion plausibility. We begin by providing a brief background about the foundational techniques (Sec. 3.1) that form the basis of our approach.
@@ 3.2 Physics-Based Motion Prior
Due to the coarse grid resolution of MPM, we refine particle-scene interactions through mesh-based collision handling. For snow, we project particles onto mesh surfaces and interpolate with nearby surface Gaussians to create natural accumulation. For rain, we implement a 3D wetness grid that tracks moisture with Gaussian smoothing and temporal decay. For sandstorms, we displace particles along surface normals with anisotropic scaling to create thin accumulations. When particles become inactive ($a_{i}^{t}=1$ to $a_{i}^{t+1}=0$), their final positions and rotations are recorded as collision states. These simulated trajectories provide physically plausible reference motion that guides our neural dynamics model. Additional collision handling details are provided in the supplementary (Sec. 8.2) and simulation parameters in Sec. 8.3.
@@ 3.3 Physics-Guided Score Distillation
Rather than treating physics simulation as fixed motion, we use it as a guidance prior to optimize a neural dynamics model through Video-SDS. For each dynamic particle from the MPM simulation, we create a corresponding dynamic Gaussian kernel $\mathcal{G}_{g}(t)=\{\mathbf{x}_{g}^{\text{init}}(t),\mathbf{R}_{g}^{\text{init}},\mathbf{A}_{g}^{\text{active}},\mathbf{A}_{g}^{\text{collided}},\mathbf{v}_{g}^{\text{init}}(t)\}$, where $\mathbf{v}_{g}^{\text{init}}(t)$ is the velocity of each gaussian that is provided along the position $\mathbf{x}_{g}^{\text{init}}(t)$ directly from the motion simulation. Each Gaussian also maintains state-dependent appearance parameters based on its physical state determined by the physics simulation: $\mathcal{A}_{g}^{\text{active}}=\{sigma_{g}^{\text{active}},\mathbf{S}_{g}^{\text{active}},\mathbf{C}_{g}^{\text{active}}\}$ for actively moving Gaussians and $\mathcal{A}_{g}^{\text{collided}}$ for Gaussians that have accumulated on static geometry. For notational clarity, we omit the explicit state subscripts and use $\mathcal{A}_{g}=\{\sigma_{g}^{\text{collided}},\mathbf{S}_{g}^{\text{collided}},\mathbf{C}_{g}^{\text{collided}}\}$ to refer to the appropriate appearance parameters based on the Gaussian’s current physical state. Appearance parameters are initialized using a large language model (LLM) that generates appropriate baseline values $\mathcal{A}_{g}^{\text{init}}$ from text prompts describing the desired weather effect. Additional details on LLM initialization are provided in the supplementary (Sec. 8.4).
Recurrent Neural Dynamics Model. To model the continuous evolution of the dynamic Gaussians, we learn a recurrent neural dynamics model that predicts dynamic Gaussian evolution over time. At each timestep $t$, the model takes as input: (1) the previous rendered state $\{\mathbf{x}_{g}(t-1),\mathbf{R}_{g}(t-1),\mathcal{A}_{g}(t-1)\}$ from the previous timestep, where $\mathbf{x}_{g}(t-1)$ is the position of Gaussian $g$ at time t1t-1, $\mathbf{R}_{g}(t-1)$ is the rotation of Gaussian $g$ and $\mathcal{A}_{g}$ represents the appearance parameters of Gaussian $g$ (opacity, scale, and color), (2) the physics velocity $\mathbf{v}_{g}^{\text{init}}(t)$ of Gaussian $g$ at time $t$ from simulation, and (3) the current timestep $t$. We implement separate MLPs for each physical state, with each MLP processing Gaussians independently. The positions, velocities, and timestep are processed through learned embedders (Fourier feature encoding for spatial inputs, sinusoidal encoding for time) before being concatenated with the current appearance parameters $\mathcal{A}_{g}(t-1)$ and fed to the appropriate state-specific MLP. The recurrence is initialized at $t=0$ using physics simulation positions $\mathbf{x}_{g}^{\text{init}}(0)$ and the LLM-initialized appearance parameters $\mathcal{A}_{g}^{\text{init}}$. Additional details on the network architecture are provided in the supplementary (Sec. 8.1).
@@
Each regularization term has independent weight parameters to provide control over physics-appearance trade-off. Additional details are provided in the supplementary (Sec. 8.1).
@@
Figure 3: Dynamic weather effects created with our physics-guided SDS framework across different scenes (input frames on left). Each column represents an increasing timestep, demonstrating temporal evolution with continuous particle emission and realistic scene interactions. Our approach enables scene-wide modifications where emitted particles interact naturally with static geometry and evolve coherently over time. The snowfall effect is presented in Fig. 1. As the rainfall effect is challenging to visualize, we provide zoomed-in regions to better illustrate the effect. Complete temporal visualizations for all effects are available in the supplementary (Sec. 6).
@@ 4.1 Qualitative Results
Fig. 1 (snowfall) and Fig. 3 (all other dynamic weather effects) demonstrate our method’s ability to create visually convincing and temporally consistent dynamic scene modifications across different environments. Our physics-guided score distillation framework enables effects that exhibit physically plausible motion while maintaining photorealistic appearance: snowflakes gently accumulate on surfaces through continuous particle emission, raindrops create wetness patterns, fog diffuses through the scene, and sandstorms envelope objects with physically-driven particle motion. Additionally, Fig. 4 showcases creative appearance variations generated by our framework through text guidance (e.g., purple snow) while maintaining physically plausible motion through our physics-guided optimization. Please refer to the supplementary (Sec. 6) for additional visualizations across different scenes, effects, and viewpoints.
@@ 4.2 Comparison To Baselines
Table 1: Quantitative Evaluation. We compare against the static 3D editing methods ClimateNeRF [18] and GaussCtrl [36], and the 4D editing method Instruct-4DGS [16]. Results are shown for different edit categories, with F+S representing results for fog and snow edits (the only effects supported by ClimateNeRF) and All representing results across all edit types.
@@
4D Instruct-4DGS $0.21$ $0.07$ $0.57$ $0.13$ $0.19$
Ours $\mathbf{0.28}$ $\mathbf{0.11}$ $\mathbf{0.89}$ $\mathbf{0.19}$ $\mathbf{0.41}$
Baselines. Our work addresses the challenge of introducing dynamic weather effects to static scenes. We compare against static 3D editing methods: ClimateNeRF [18], which adds static weather effects to NeRFs, and GaussCtrl [36], a method for text-guided editing of 3D Gaussian Splatting scenes. We also compare against Instruct-4DGS [16], a recent 4D editing method that applies text-guided edits to dynamic Gaussian scenes via score distillation. Since Instruct-4DGS assumes an existing dynamic scene, we initialize its deformation field to zero and train it jointly, allowing the method to learn motion dynamics from diffusion guidance (additional details in the supplementary, Sec. 7). RainyGS [4] addresses rain synthesis through physics-based particle systems with rain-specific simulation and rasterization techniques, however, it is not directly comparable as it does not generalize beyond rain effects and their code is not publicly available.
@@
Comparisons. A qualitative comparison to static 3D editing methods is shown in Fig. 5. We show our results at two different timesteps: one at an early stage of the simulation (Small $t$) and one at the end (Large $t$). In the snowfall example in the first row, the static methods cannot capture the temporal dynamics of continuous particle emission and accumulation, yielding only surface modifications (ClimateNeRF) or appearance changes (GaussCtrl). In the second row, for the fog example, the effect is either unnoticeable (ClimateNeRF) or unrealistic (GaussCtrl). Our physics-guided framework produces effects that exhibit both physically plausible temporal evolution and photorealistic appearance. Additional comparisons across different scenes, effects, and viewpoints are provided in the supplementary (Sec. 7, Sec. 9).
In Tab. 1 we show a quantitative comparison to to these baselines. For the static methods, we create moving camera videos from their results to enable fair comparison using video-based metrics. ClimateNeRF only performs modifications for the fog and snow effects, thus we show results over these effects (F+S) and also over all weather effects. As illustrated in the table, our method achieves higher scores across both video and image-based metrics, demonstrating that our approach produces high-quality effects even in static frames, while the temporal dimension, enabled by continuous particle emission and physics-guided optimization, further distinguishes our method. We additionally compare against Instruct-4DGS in Tab. 1 (4D). Without a physics-based motion prior, Instruct-4DGS produces incoherent results with physically implausible motion, resulting in significantly lower scores across all metrics (qualitative comparison in the supplementary, Fig. 7).
@@ 4.3 Ablations
Fig. 6 and Tab. 2 illustrate that each ablated component leads to a significant degradation in quality. The w/o CH ablation demonstrates that without proper collision handling, dynamic Gaussians fail to interact convincingly with static scene geometry, with snowflakes floating unnaturally above surfaces instead of accumulating realistically. While the w/o App variant achieves physically plausible motion, it produces suboptimal visual appearance due to simplified LLM-initialized parameters. The w/o Motion variant produces incoherent and noisy results, producing significantly lower results than other variants. These results demonstrate that video diffusion models alone cannot reliably generate physically plausible motion for complex multi-particle scenarios and are not designed to introduce new Gaussians. The w/o PG variant demonstrates that fixed motion prevents effective Video-SDS optimization. Without the joint optimization enabled by our physics-guidance prior, the model cannot learn the motion refinements required for the diffusion model to satisfy the text prompt. Additional ablation results are provided in the supplementary (Sec. 7).
@@ 5 Conclusion
Acknowledgements.
This research was supported by The Israel Science Foundation (grant No. 2416/25).
Supplementary Material
6 Interactive Visualizations
We refer readers to the interactive visualizations at https://galfiebelman.github.io/let-it-snow/supp/index.html for full temporal sequences of dynamic weather effects, comparisons to 4D editing baselines, and comprehensive ablation studies.
7 Additional Ablations and Comparisons
@@
As shown in Fig. 7 (quantitative results in Tab. 1 of the main paper), Instruct-4DGS produces incoherent results with physically implausible motion patterns. Without a physics-based motion prior to guide the optimization, the method struggles to learn realistic dynamics for complex multi-particle weather phenomena, resulting in unnatural motion and visual artifacts.
@@
Physics Prior Robustness. We evaluate the robustness of our framework to perturbations in key simulation parameters. We perturb four parameters by $\pm$50%: emission rate, initial velocity, Young’s modulus, and movement threshold $\delta$. Experiments are conducted on all scenes across snowfall, rainfall, and sandstorm effects. As shown in Tab. 6, our method maintains stable performance across all perturbations and motion remains stable with average trajectory drift (Eq. 9) of $0.28\pm 0.03$ compared to a baseline drift of $0.26$ on these three effects.
Table 6: Physics Prior Robustness. Performance under $\pm$50% perturbation of key simulation parameters (emission rate, initial velocity, Young’s modulus, movement threshold), averaged across all scenes and three effects (snowfall, rainfall, sandstorm). Our method maintains stable performance across all perturbations.
Setting $\text{CLIP}_{Sim}\uparrow$ $\text{CLIP}_{Dir}\uparrow$ $\text{VQAScore}\uparrow$ $\text{ViCLIP-T}\uparrow$ $\text{VE-Bench}\uparrow$
$\pm$50% Perturbation $0.27{\pm}0.02$ $0.10{\pm}0.01$ $0.87{\pm}0.04$ $0.17{\pm}0.01$ $0.38{\pm}0.03$
Default (Ours) $\mathbf{0.28}$ $\mathbf{0.11}$ $\mathbf{0.88}$ $\mathbf{0.19}$ $\mathbf{0.39}$
Regularization Weight Sensitivity. We vary each regularization weight independently by $\times 0.1$ and $\times 10$ while keeping others at their default values. Experiments are conducted on all scenes across snowfall and sandstorm effects. As shown in Tab. 7, physics regularization weights ($\lambda_{\text{xyz}}$, $\lambda_{\text{vel}}$, $\lambda_{\text{rot}}$) are robust to order-of-magnitude changes, with metrics remaining within 0.01-0.02 of default values. Appearance regularization weights are more sensitive when reduced by $10\times$ but degrade gracefully when increased, confirming a stable operating range around our defaults.
Table 7: Regularization Weight Sensitivity. Each weight is varied independently by $\times 0.1$ and $\times 10$ on all scenes across snowfall and sandstorm effects. Physics weights are robust to order-of-magnitude changes. Appearance weights are sensitive when reduced but degrade gracefully when increased.
Weight Value $\text{CLIP}_{Sim}\uparrow$ $\text{CLIP}_{Dir}\uparrow$ $\text{VQAScore}\uparrow$ $\text{ViCLIP-T}\uparrow$ $\text{VE-Bench}\uparrow$
$\lambda_{\text{xyz}}$ (0.1) $\times 0.1$ 0.26 0.09 0.84 0.16 0.35
$\times 10$ 0.27 0.10 0.86 0.17 0.37
$\lambda_{\text{vel}}$ (0.1) $\times 0.1$ 0.26 0.10 0.85 0.16 0.37
$\times 10$ 0.27 0.10 0.86 0.17 0.39
$\lambda_{\text{rot}}$ (0.1) $\times 0.1$ 0.28 $\mathbf{0.11}$ 0.88 $\mathbf{0.19}$ 0.39
$\times 10$ 0.28 $\mathbf{0.11}$ 0.88 $\mathbf{0.19}$ 0.41
$\lambda^{\text{active}}_{\sigma}$ (1.0) $\times 0.1$ 0.24 0.08 0.79 0.15 0.32
$\times 10$ 0.27 0.10 0.86 0.18 0.39
$\lambda^{\text{active}}_{\mathbf{S}}$ (1.0) $\times 0.1$ 0.19 0.04 0.52 0.09 0.19
$\times 10$ 0.26 0.10 0.83 0.17 0.37
$\lambda^{\text{active}}_{\mathbf{C}}$ (35.0) $\times 0.1$ 0.20 0.06 0.57 0.11 0.23
$\times 10$ 0.26 0.10 0.84 0.18 0.37
$\lambda^{\text{collided}}_{\sigma}$ (35.0) $\times 0.1$ 0.22 0.07 0.74 0.13 0.30
$\times 10$ 0.27 0.10 0.85 0.18 0.38
$\lambda^{\text{collided}}_{\mathbf{S}}$ (35.0) $\times 0.1$ 0.17 0.03 0.47 0.07 0.14
$\times 10$ 0.26 0.10 0.84 0.18 0.38
$\lambda^{\text{collided}}_{\mathbf{C}}$ (35.0) $\times 0.1$ 0.21 0.06 0.61 0.12 0.26
$\times 10$ 0.27 0.10 0.85 0.18 0.37
Default (Ours) $\mathbf{0.29}$ $\mathbf{0.11}$ $\mathbf{0.89}$ $\mathbf{0.19}$ $\mathbf{0.42}$
Scheduler Comparison. We compare our SDS-adaptive scaling against three fixed regularization weight decay schedules on all scenes across snowfall and sandstorm effects: (1) linear decay ($\lambda$: $10^{6}\rightarrow 10^{2}$ over 1000 iterations), (2) exponential decay ($\lambda{=}10^{6}\cdot 0.995^{\text{iter}}$), and (3) cosine annealing ($\lambda$: $10^{6}\rightarrow 10^{2}$ following a cosine curve). As shown in Tab. 8, our SDS-adaptive scaling outperforms all fixed schedules. Fixed schedules fail because Video-SDS loss magnitude varies unpredictably during training, while our adaptive scaling strengthens physics guidance when SDS is uncertain and relaxes it when the model is confident.
Table 8: Scheduler Comparison. Comparison of fixed regularization weight decay schedules versus our SDS-adaptive scaling on all scenes across snowfall and sandstorm effects. Our adaptive approach outperforms all fixed schedules.
Schedule $\text{CLIP}_{Sim}\uparrow$ $\text{CLIP}_{Dir}\uparrow$ $\text{VQAScore}\uparrow$ $\text{ViCLIP-T}\uparrow$ $\text{VE-Bench}\uparrow$
Linear Decay 0.25 0.09 0.82 0.15 0.35
Exponential Decay 0.24 0.08 0.80 0.14 0.34
Cosine Annealing 0.26 0.10 0.84 0.17 0.38
Ours (SDS-Adaptive) $\mathbf{0.29}$ $\mathbf{0.11}$ $\mathbf{0.89}$ $\mathbf{0.19}$ $\mathbf{0.42}$
Per-Particle Appearance and Rotation Ablation. We ablate the necessity of per-particle appearance variation and angular velocity prediction. We evaluate two variants on all scenes across snowfall and sandstorm effects: (1) fixed appearance, where all active particles share one optimized appearance and all collided particles share another, and (2) fixed rotation, where the angular velocity prediction is disabled. The variance of learned parameters across particles, averaged over timesteps, scenes, and effects, confirms meaningful per-particle variation for appearance: $\text{Var}(\sigma)_{\text{active}}{=}0.12$, $\text{Var}(\mathbf{S})_{\text{active}}{=}0.21$, $\text{Var}(\mathbf{C})_{\text{active}}{=}0.05$, $\text{Var}(\sigma)_{\text{collided}}{=}0.01$, $\text{Var}(\mathbf{S})_{\text{collided}}{=}0.03$, $\text{Var}(\mathbf{C})_{\text{collided}}{=}0.005$, while angular velocity variance is near-zero ($\text{Var}(\omega){\approx}0$).
Tab. 9 demonstrates that removing per-particle appearance clearly degrades performance across all metrics, confirming its importance. Fixed rotation shows no degradation, consistent with the near-zero angular velocity variance ($\text{Var}(\omega){\approx}0$). We retain this capacity for generality and for effects where rotation may be more prominent.
Table 9: Per-Particle Appearance and Rotation Ablation. Removing per-particle appearance degrades all metrics. Removing angular velocity shows no degradation, consistent with near-zero learned rotation variance for weather effects.
Variant $\text{CLIP}_{Sim}\uparrow$ $\text{CLIP}_{Dir}\uparrow$ $\text{VQAScore}\uparrow$ $\text{ViCLIP-T}\uparrow$ $\text{VE-Bench}\uparrow$
Fixed Appearance 0.26 0.09 0.83 0.17 0.36
Fixed Rotation $\mathbf{0.29}$ $\mathbf{0.11}$ 0.88 $\mathbf{0.19}$ $\mathbf{0.43}$
Ours $\mathbf{0.29}$ $\mathbf{0.11}$ $\mathbf{0.89}$ $\mathbf{0.19}$ 0.42
SDS Variant Comparison. We evaluate two improved score distillation variants to assess compatibility with our framework, Noise-Free Score Distillation (NFSD) [13], which decomposes the score into interpretable components and removes an undesired noise term to reduce over-smoothing, and Bridge-SDS [22], which formulates score distillation as optimal transport between source and target distributions to reduce oversaturation artifacts. We replace our vanilla SDS objective with each variant and evaluate on all scenes across snowfall and sandstorm effects, keeping all other components identical.
As shown in Tab. 10, both improved variants yield higher scores across all metrics, with Bridge-SDS achieving the strongest results. While our physics-guided optimization drives the primary quality improvements over baselines (Tab. 1), these results confirm that orthogonal advances in score distillation provide additional gains when integrated into our framework.
Table 10: SDS Variant Comparison. Our framework is compatible with improved score distillation methods. Both NFSD [13] and Bridge-SDS [22] improve over vanilla SDS when integrated into our physics-guided optimization.
SDS Variant $\text{CLIP}_{Sim}\uparrow$ $\text{CLIP}_{Dir}\uparrow$ $\text{VQAScore}\uparrow$ $\text{ViCLIP-T}\uparrow$ $\text{VE-Bench}\uparrow$
Vanilla SDS 0.29 0.11 0.89 0.19 0.42
NFSD 0.30 0.11 0.91 0.20 0.44
Bridge-SDS $\mathbf{0.32}$ $\mathbf{0.12}$ $\mathbf{0.93}$ $\mathbf{0.22}$ $\mathbf{0.45}$
Comparison to Manual VFX. To contextualize the quality-automation tradeoff, we created an equivalent snowfall effect in Blender [3] on the Garden scene. As a novice user, manual setup in Blender required approximately 10 hours (particle system configuration, material setup, lighting adjustment, rendering), compared to our automated pipeline of approximately 1 hour. A visual comparison is provided in Fig. 11. Our results are comparable or better than the manually created Blender effect, suggesting that little to no realism is lost for the sake of automation while requiring significantly less time and expertise.
xxxxxxxxBlender
Refer to caption Refer to caption
xxxxxxxxOurs
Refer to caption Refer to caption
Figure 11: Comparison to Manual VFX. Two consecutive frames of a snowfall effect on the Garden scene created manually in Blender [3] ($\sim$}10 hours, top) versus our automated pipeline ($\sim$}1 hour, bottom). Our automated pipeline produces results of comparable visual quality while requiring an order of magnitude less time and no manual expertise.
8 Additional Details
8.1 Implementation Details